Find the value of a³+2a²b+b³ divided by ab(a+3b) when a/b=√2-1


1
Expert's answer
2021-05-17T04:02:02-0400

We have given that,


"\dfrac{a}{b} = \sqrt{2}-1"


"a = b(\sqrt{2}-1)"

We have to find the value of "\dfrac{a^3+2a^b+b^3}{ab(a+3b)}"


Putting the value of a in the above equation,


We get,


"\dfrac{b^3(\sqrt{2}-1)^3+2b^3(\sqrt{2}-1)^2+b^3}{b^3(\sqrt{2}-1)^2+3b^3(\sqrt{2}-1)}"


"= \dfrac{(\sqrt{2}-1)^3+(\sqrt{2}-1)^2+1}{(\sqrt{2}-1)^2+(\sqrt{2}-1)}"


After solving we get,


"\dfrac{3 \sqrt{2}-1}{2-\sqrt{2}}"


"= \dfrac{3}{\sqrt{2}}"